# S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study

### v1

@@ Abstract
In the resource-constrained world of the digital landscape, lightweight cryptography plays a critical role in safeguarding information and ensuring the security of various systems, devices, and communication channels. Its efficient and resource-friendly nature makes it the ideal solution for applications where computational power is limited. In response to the growing need for platform-specific implementations, NIST issued a call for standardization of lightweight cryptography algorithms in 2018. Ascon emerged as the winner of this competition. NIST initally established general evaluation criteria for a standard lightweight scheme including security strength, mitigation against side-channel and fault-injection attacks, and implementation efficiency. To verify the security claims, evaluating the individual components used in any cryptographic algorithm is a crucial step. Substitution box (sbox) is one of the fundamental components used to ensure Shannon’s property of confusion in cryptographic primitives. The quality of a substitution box (S-box) significantly impacts the overall security of a cryptographic primitive. This paper analyzes the S-boxes of six finalists in the NIST Lightweight Cryptography (LWC) standardization process. We evaluate them based on well-established cryptographic properties. Our analysis explores how these properties influence the S-boxes’ resistance against known cryptanalytic attacks and potential implementation-specific vulnerabilities, thus reflecting on their compliance with NIST’s security requirements.
@@
Sbox, Lightweight Cryptography, NIST, Substitution Layer
volume: 185issue: 1
S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study
@@ 1 Introduction
The growing need for compact, resource-constrained devices performing a variety of functions including but not limited to sensing, identification, and machine control has driven the development of lightweight cryptographic designs and algorithms. The highly evolving yet resource constrained environment raised countless security concerns. As a result, using traditional cryptographic primitives on these platforms while striking an optimum trade-off between compact implementation and security strength became difficult. In order to effectively resolve this issue, NIST lay the foundation for lightweight cryptography standardization in $2017$ [1],[2].
@@
Stemming from the design principles of symmetric cryptography, Lightweight cryptography aims to provide security for Internet of Things (IoT) devices namely sensors, RFID tags, medical implants, smart cards, etc. Design choices such as small block and key size, simple round structure and key schedule and nominal implementation make lightweight ciphers the right choice to ensure a sufficient level of security using less power, computing and memory resources.
@@
As a result of NIST’s call, a total of $57$ proposals were submitted by research teams across the globe [3]. After two rounds of rigorous evaluation process, as described in NIST IR $8369$ [4], $10$ candidates were shortlisted as finalists of the competition, see table 1 In the third and final round, candidates were evaluated on the basis of the claims made in terms of security strength against known attacks ( 112 bits or more); performance on both hardware and software environment; resilience against side channel and fault injection based attacks; and intellectual property rights [5].
@@
As a result, in February $2023$, NIST crowned ASCON family as the new lightweight cryptography standard [6]. The most important evaluation criteria for cryptographic algorithms is measuring their security strength against known attacks such as differential and linear cryptanalysis (and their variants), as well as implementation based attacks. Designers and cryptanalysts across the globe tried their level best to suffice this requirement. As a result, the competition finalists received a noticeable amount of third party security analysis, selective ones highlighted as follows:-
  • Cihangir T. performed key recovery using differential-linear attack on $4$ $5$ rounds of ASCON with a time complexity of $2^{15}$ $2^{33.1}$ [7]. In [8], used undisturbed bits in sbox to perform impossible, truncated and improbable differential attacks on $4$ and $5$ round ASCON. Yanbin L. et al. performed key recovery attack on ASCON $7/12$ round initialization in the nonce misuse scenario with a time complexity of $2^{97}$. They also suggested a forgery attack on $4$, $5$ and $6$ round finalization with timing complexity $2^{9}$, $2^{17}$ $2^{33}$[9]. Maria et al. performed cube based attack on $5$ $6$ rounds with time complexity $2^{35}$ $2^{66}$ and differential-linear key recovery attack on $4$ $5$ rounds of ASCON with a complexity of $2^{18}$ $2^{36}$ [10]. Duarte-Sanchez et al. presented a combined attack using a round reduction hardware trojan to reduce the number of rounds in Ascon to 5 in the initialization phase, and then perform a cube key recovery attack with a timing complexity of $2^{24}$ Further analysis of ASCON includes [11], [12, 13, 14, 15] etc.
  • F.X. et al. show that single trace attacks against leakage resistant PRF used in ISAP can be better mitigated by combining rekeying and shuffling instead of rekeying and masking as suggested in [16]. They also concluded that ISAP is better suited to ensure side channel security on parallelized architectures as compared to low end devices [17].
  • In [18] Yaxin et al. use MILP to find linear characteristics of GIFT-$128$ and then launch a linear cryptanalysis attack on $20$ rounds in order to guess $65$ bits of key. Cao et al. make use of Sun et al. MILP based tool [19] to search for differential characteristics of $7$ $10$ rounds of GIFT-$128$ having a probability of $2^{-15.83}$ $2^{-72.66}$ [20]. Andreas et al. proposed a tool based on bit division property to search for an integral distinguisher on the $11$ rounds of GIFT-$128$ [21]. Wentao et al. used Matsui’s technique to perform related key differential attack on the GIFT-$128$ and locate the best differential trail for upto $14$ rounds [22]. Other attacks include [23, 24, 25, 26] etc.
  • A. Jana and G. Paul proposed two fault models to perform forgery attack on Photon-Beetle in order to first recover the state and then $128$-bit master key. In random fault model, $2^{37.15}$ fault queries, and in known fault model $2^{11.05}$ fault queries are required to recover the entire key [27]. Bart et al. noticed an irregularity in the claimed security bounds leading to a possibility of key recovery attack with $2^{124}$ queries [28]. Other attacks include [29, 30, 31, 32] etc.
  • Haibou Z. et al. proposed interpolation attack on $8$ rounds of Delirium, using $2^{23.1}$ linear equations to perform key recovery with attack complexity $2^{70}$ [33]. Priyanka J. et al. proposed single event fault analysis attack on Dumbo to perform full key recovery using $85$ to $250$ cipher text pairs [34]. A.H. Meraneh et al. theoretically prove the possibility of blind side channel attack on the LFSR used in Elephant. As the internal state of the LFSR is dependant on the secret key. Hence in most use cases this security compromise could lead to secret key being recovered in less than two days [35]. Further attacks include [36, 37].
  • Ahmed A. et al. present an impossible differential attack on $22$ rounds of Skinny-$128$ (underlying cipher in Romulus) [38]. Kazuhiko et al. proposed matching and authenticity attack on Romulus-M with tight security bounds [39]. J. Hua et al. devised a constrained framework to find Meet in the Middle (MITM) characteristics for performing single key recovery attack on $23$ rounds of Skinny-$128$ with time and data complexity of $2^{376}$ $2^{104}$[40]. Other attacks include [41, 42, 43, 44] etc.
Table 1: Methods of ensuring integrity of digital evidence
@@
 [^1] Current paper is an extended version of our previous paper [45]
Shannon’s confusion in any symmetric cipher is mainly dependent on the non-linear component used. The most commonly used component to achieve non-linearity in symmetric cryptography is substitution box (sbox). Analysing the cryptographic profile of an sbox plays a vital role in determining the overall strength of a cryptosystem. Literature reports nearly $28$ cryptographic properties [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] that can be used as a tool to estimate the quality of an sbox. Some of the attacks mentioned earlier exploit specific sbox properties in order to devise attack characteristics for a particular cipher. For example, in [34] authors point out that the value of correlation immunity of ASCON sbox makes it susceptible to subset cryptanalysis. However, an entire picture of the cryptographic properties of finalist sboxes is not present in literature and our aim is to fill that gap. Section 2 of this paper elaborates the specifications of six finalists of NIST LWC. Section 3 discusses the fundamental properties of sboxes and classifies them on the basis of relevant attack category. Secondly, upper and lower bounds on these properties are highlighted along with their behaviour under affine equivalence. In Section 4 the results of sbox analysis are discussed in detail. Conclusion is presented in Section 5.
@@ 2 Overview of Finalists
This section provides a brief introduction to each of the $6$ finalists from NIST lightweight standardization that make use of sbox as one of the design components.
@@ 2.1 ASCON
It is a cryptographic cipher suite presented by Maria et al. [62]. ASCON comprises of three AEAD variants: $128$, $128$ ($a$ and $pq$), and two hashing algorithms: Hash and Hash-$a$. Its main functionality comes from the use of a $320$ bit permutation which is combined with different constants and number of rounds, depending upon the variant. In addition, to achieve non-linearity in the encryption algorithm a $5\times 5$ affine equivalent sbox of $\chi$ mapping from Keccak permutation is used. The ASCON family has been selected by NIST for lightweight standardization because of its security strength and flexibility to be used in multiple applications while bearing a minimum implementation cost [6].
@@ 2.2 ISAP
ISAP was proposed by S. Mangard et al. [63], it is a nonce based AEAD cryptographic scheme designed to resist against implementation based attacks at an optimized implementation cost. ISAP uses $400$ bit Keccak and $320$ bit ASCON permutations in its design alongside $5\times 5$ bit ASCON sbox as a non-linear component.
@@ 2.3 GIFT-COFB
Proposed by A. Chakraborti. et al.[64]. GIFT-COFB is an encryption scheme that uses combined feedback mode with GIFT-$128$. GIFT-COFB employs $40$ rounds of encryption where the input is: an arbitrary length associated data, an input message of variable length, a nonce of $128$ bits, and a $128$ bit encryption key. Whereas the output is a $128$ bit tag and ciphertext of same length as the input message. A $4\times 4$ bit sbox is used as the confusion component.
@@ 2.4 Photon-Beetle
Photon-Beetle is an AEAD based hashing and encryption scheme presented by J. Guo et al.[65]. The encryption routine comprises of $12$ rounds consisting of $4$ separate stages i.e. XOR constant, substitute cells, shift rows and serial mix columns. Photon-Beetle (AEAD) uses the $4$ $4$ bit sbox of Present block cipher in order to achieve non-linearity in its design.
@@ 2.5 Elephant
C. Dobarunig et al. [66], proposed Elephant as a permutation based AEAD algorithm. It has $3$ variants based on Spongent and Keccak permutations: Delirium, Jumbo and Dumbo. These encryption variants take $128$ bit key and $96$ bit nonce, however the block sizes are $160$, $176$, $200$. A $4\times 4$ sbox as the non-linear part. In the $80$ rounds of encryption there are $160$ differentially active sboxes.
@@ 2.6 Romulus
T. Iwata. et al. [67], proposed Romulus as a nonce based AEAD encryption and hashing scheme that uses Skinny block cipher as underlying encryption algorithm. It operates in $4$ modes of encryption and hashing: nonce misuse based Romulus-M, nonce based Romulus-N, leakage resistant based Romulus-T and hashing based Romulus-H. The encryption algorithm takes $128$ bits of key and nonce, and runs $40$ rounds of encryption. Skinny-$128$ tweakable block cipher [68] uses $8\times 8$ sbox to achieve non-linearity in its design.
Table 2: Classification of s-box properties as per
relevance to existing cryptanalysis techniques
@@
1 Generic
Balancedness
Permutation
Order of Permutation
Fixed Points (FP)
Opposite Fixed Points (OFP)
Bit Independence Criteria (BIC)
Strict Avalanche Criteria (SAC)
Auto Correlation Table (ACT) [69]
Absolute Indicator (AI) [70]
Sum of Square Indicators (SSI) [70]
2 Linear Cryptanalysis
Nonlinearity (NL) [50, 51]
Linear Approximation Table (LAT) [52, 71]
Linear Approximation Probability (LAP)
Linear Branch Number (LBN)
Linear Structures (LS) [53, 54]
Correlation Immunity (CI) [72]
3 Differential Cryptanalysis
Differential Distribution Table (DDT) [52]
Differential Uniformity (DU) [50, 51]
Differential Branch Number (DBN)
Propagation Characteristics (PC) [73]
Undisturbed Bits [56]
4 Boomerang Cryptanalysis
Boomerang Connectivity Table (BCT)
Boomerang Uniformity (BU) [55]
4 Differential - Linear Cryptanalysis
Linear Structures (LS) [53, 54]
Differential Linear Connectivity Table (DLCT) [57]
Differential Linear Uniformity (DLU) [58]
5 Algebraic Cryptanalysis
Algebraic Degree (AD) [59]
Interpolation Polynomial (IP)
6 Side Channel Analysis
DPA-SNR[60]
Transparency Order (TO) [61]
Table 3: Theoretical Bounds on the Cryptographic Properties of Sboxes
S# Properties
Bounds

### v3

@@ Abstract
Lightweight cryptography was primarily inspired by the design criteria of symmetric cryptography. It plays a vital role in ensuring the security, privacy, and reliability of microelectronic devices without compromising the overall functionality and efficiency. However, the increasingly platform specific design requirements prompted the development of a standard lightweight algorithm. In 2017, NIST put forward security requirements for a standard lightweight scheme — security strength of at least 112 bits against known cryptanalysis attacks, mitigation against side channel and fault injection attacks, and implementation efficiency. After three rounds of review, ASCON was crowned as the winner of the competition. Evaluating the individual components used in any cryptographic algorithm is an important step in the verification of security claims. A fundamental component used to ensure Shannon’s property of confusion in cryptographic primitives is an S-box. Hence, the quality of an S-box is a significant contributing factor in the security strength of a cipher. In this paper, we evaluate the S-boxes of 6 NIST LWC competition finalists based on well-known cryptographic properties, and comment on how the results reflect upon NIST security requirements. Our findings have revealed that these S-boxes do not comply with the basic notions of avalanche, making it vulnerable to high-order sophisticated cryptanalysis.
@@
NIST, Lightweight Cryptography, Substitution layer, Cryptanalysis, S-box
@@ 1 Introduction
Application specific computing has become an emerging trend in modern information technology due to the wide usability of smart devices. The highly evolving yet resource constrained environment enumerates to countless security concerns. It is difficult to use traditional cryptographic primitives on these platforms while striking an optimal trade-off between compact implementation and security strength [1]. In order to effectively resolve this issue, NIST laid the foundation for lightweight cryptography standardization in $2017$ [2].
@@
Lightweight cryptography aims to provide security for Internet of Things (IoT) devices, namely sensors, RFID tags, medical implants, smart cards, etc. Design choices such as small block and key size, simple round structure and key schedule, and nominal implementation make lightweight ciphers the right choice to ensure a sufficient level of security using less power, computing, and memory resources.
@@
As a result of NIST’s call, a total of $57$ proposals were submitted by research teams across the globe. After two rounds of the rigorous evaluation process, as described in NIST IR $8369$ [3], $10$ candidates were shortlisted as finalists of the competition (see Table 1 [4]). In the third and final round, candidates were evaluated on the basis of claims made in terms of security strength against known attacks ($112$ bits or more); performance in both hardware and software environments; resilience against side channel and fault injection attacks; and intellectual property rights [5].
@@
In February $2023$, NIST declared the ASCON family as the new lightweight cryptography standard [6]. The most important evaluation criteria for cryptographic algorithms is measuring their security strength against known attacks such as differential and linear cryptanalysis (and their variants), as well as implementation based attacks. Designers and cryptanalysts across the globe tried their level best to suffice this requirement. As a result, the competition finalists received a noticeable amount of third-party security analysis [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], a few highlighted in Table 2:-
Table 1: Classification of NIST LWC Finalists based on Presence/Absence of S-box
@@
Table 2: Summary of Third-Party Security Analysis on NIST LWC Finalists
Cipher Attack Type Rounds Complexity Reference
ASCON Differential-Linear 44 55 Time $2^{15}$, $2^{18}$ Time $2^{33.1}$, $2^{36}$ [7] [9]
Impossible Differential Truncated Differential $4,5$ $4,5$ Data $2^{2}$, $2^{109}$ Data $2^{2}$, $2^{109}$ [45]
Cube Attack 55 66 Time $2^{35}$ Time $2^{66}$ [9]
Key Recovery (Nonce Misuse) Init ($7/12$) $8,9$ Time $2^{97}$, $2^{101}$, $2^{123.92}$ [8] [46]
Forgery Attack Final (46)(4-6) Time $2^{9}$, $2^{17}$ & $2^{33}$ [8]
Round Reduction ++ Cube Init $(5)$ Time $2^{24}$ [10]
CPA using Multi-bit Selection Function Full 22-bit selection function =7200=7200 traces $3$-bit selection function =8200=8200 traces [47]
Differential Fault Attack Full Fault queries $2^{10}$ [48]
Collision Attack 22 Time $2^{98}$ [49]
Free-start Collison Attack 33 Time $2^{10}$ [50]
GIFT-128 / COFB MILP-Based Linear Attack 2020 $65$-bit key recovery [18]
Differential Key Recovery Attack $23$ Time $2^{120}$, Data $2^{86}$ [51]
Integral Distinguisher using Division Property $11$ - [20]
Differential Attack $26$ Time $2^{123.245}$ [21]
Photon-Beetle Key Recovery Irregularity Full Data $2^{50}$ bytes [28]
Differential Fault Attack Full Time $2^{16}$, Fault queries $2^{37.15}$ [52]
Committing Attack Initial State No queries [53]
Pre-image Attack $3.5$ Time = $2^{112}$, Data $2^{65}$ [54]
Elephant Interpolation Attack 88 Time $2^{70}$ [33]
Fault Analysis (Dumbo) Full 8525085-250 ciphertexts [34]
Blind SCA on LFSR Full <2<2 days (practical) [35]
Romulus Impossible Differential $22$ Time $2^{373.48}$ , Data $2^{92.22}$ [38]
Matching / Authenticity Attack Full - [39]
Meet-in-the-Middle (MITM) $23$ Time $2^{376}$ , Data $2^{104}$ [40]
Impossible Boomerang $27$ $28$ Time $2^{337}$, Data $2^{131.3}$ Time $2^{382.8}$, Data $2^{130.26}$ [55]
Differential MITM 25 Time $2^{366}$, Data $2^{122.3}$ [56]
Impossible Differential MITM 2121 $22$ $23$ $27$ $28$ Time $2^{344.33}$, Data $2^{122.89}$ Time $2^{378.22}$, Data $2^{121.50}$ Time $2^{378.22}$, Data $2^{121.50}$ Time = $2^{361.06}$, Data = $2^{124.99}$ Time $2^{378.22}$, Data $2^{129.50}$ [57]
Shannon’s property confusion in any symmetric cipher is mainly dependent on the non-linear component used. The most commonly used component to achieve non-linearity in symmetric cryptography is a substitution box (S-box). Analysing the cryptographic profile of an S-box plays a vital role in determining the overall strength of a cryptosystem. The literature reports nearly $24$ cryptographic properties [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] that can be used as a tool to estimate the quality of an S box. Some of the attacks mentioned earlier exploit specific S-box properties in order to devise practical attack characteristics for a particular cipher. This paper aims to provide a complete picture of the cryptographic properties of S-boxes, their relevance to known attacks, and related security limits  [^1] Current paper is an extended version of our previous paper [74]. Our primary goal is to help researchers better understand and evaluate the security profile of S-boxes in a principle manner.
Section 2 of this paper elaborates the specifications of six finalists of NIST LWC. Section 3 discusses the fundamental properties of S boxes and classifies them according to the relevant attack category. Secondly, the upper and lower bounds of these properties are highlighted along with their behavior under affine equivalence. In Section 4, the results of the S-box analysis are discussed in detail. In Section 5, we generate a new 44-bit S-box with optimal cryptographic properties using the PEIGEN S-box generation tool [75], and draw a comparison of the properties of this S-box with the LWC finalists having 44-bit S-boxes. Section 6 discusses and highlights the scope and limitations of this paper. Finally, we conclude our findings in Section 7.
@@ 2 Overview of Finalists
This section provides a brief introduction to each of the $6$ finalists from NIST lightweight standardization that make use of S-box as one of the design components.
@@ 2.1 ASCON
It is a cryptographic cipher suite presented by Maria et al. [76]. ASCON comprises of three authenticated encryption and associated data (AEAD) variants: $128$, $128$ ($a$ and $pq$), and two hashing algorithms: Hash and Hash-$a$. Its main functionality comes from the use of a $320$ bit permutation, which is combined with different constants and number of rounds, depending upon the variant. In addition, to achieve non-linearity in the encryption algorithm a $5\times 5$ affine equivalent S-box of $\chi$ mapping from Keccak permutation is used. The ASCON family has been selected by NIST for lightweight standardization because of its security strength and flexibility to be used in multiple applications while bearing a minimum implementation cost [6].
@@ 2.2 ISAP
ISAP was proposed by S. Mangard et al. [77], it is a nonce based AEAD cryptographic scheme designed to resist against implementation based attacks at an optimized implementation cost. ISAP uses $400$ bit Keccak and $320$ bit ASCON permutations in its design alongside $5\times 5$ bit ASCON S-box as a non-linear component.
@@ 2.3 GIFT-COFB
Proposed by A. Chakraborti. et al.[78]. GIFT-COFB is an encryption scheme that uses combined feedback mode with GIFT-$128$. GIFT-COFB employs $40$ rounds of encryption where the input is: an arbitrary length associated data, an input message of variable length, a nonce of $128$ bits, and a $128$ bit encryption key. Whereas the output is a $128$ bit tag and ciphertext of same length as the input message. A $4\times 4$ bit S-box is used as the confusion component.
@@ 2.4 Photon-Beetle
Photon-Beetle is an AEAD based hashing and encryption scheme presented by J. Guo et al. [79]. The encryption routine comprises of $12$ rounds consisting of $4$ separate stages i.e. XOR constant, substitute cells, shift rows and serial mix columns. Photon-Beetle (AEAD) uses the $4$ × $4$ bit S-box of Present block cipher in order to achieve non-linearity in its design.
@@ 2.5 Elephant
C. Dobarunig et al. [80], proposed Elephant as a permutation based AEAD algorithm. It has $3$ variants based on Spongent and Keccak permutations: Delirium, Jumbo and Dumbo. These encryption variants take $128$ bit key and $96$ bit nonce, however the block sizes are $160$, $176$, $200$. A $4\times 4$ S-box as the non-linear part. In the $80$ rounds of encryption there are $160$ differentially active S-boxes.
@@ 2.6 Romulus
T. Iwata. et al. [81], proposed Romulus as a nonce based AEAD encryption and hashing scheme that uses Skinny block cipher as underlying encryption algorithm. It operates in $4$ modes of encryption and hashing: nonce misuse based Romulus-M, nonce based Romulus-N, leakage resistant based Romulus-T and hashing based Romulus-H. The encryption algorithm takes $128$ bits of key and nonce, and runs $40$ rounds of encryption. Skinny-$128$ tweakable block cipher [82] uses $8\times 8$ S-box to achieve non-linearity in its design.
Table 3: Classification of S-box properties as per relevance to existing cryptanalysis techniques
@@
1 Generic Balancedness; Permutation; Order of Permutation; Fixed Points (FP); Opposite Fixed Points (OFP); Bit Independence Criteria (BIC); Strict Avalanche Criteria (SAC); Auto Correlation Table (ACT) [83]; Absolute Indicator (AI) [84]; Sum of Square Indicators (SSI) [84]
2 Linear Cryptanalysis Nonlinearity (NL) [62, 63]; Linear Approximation Table (LAT) [64, 85]; Linear Approximation Probability (LAP); Linear Branch Number (LBN); Linear Structures (LS) [65, 66]; Correlation Immunity (CI) [86]
3 Differential Cryptanalysis Differential Distribution Table (DDT) [64]; Differential Uniformity (DU) [62, 63]; Differential Branch Number (DBN); Propagation Characteristics (PC) [87]; Undisturbed Bits [68]
4 Boomerang Cryptanalysis Boomerang Connectivity Table (BCT); Boomerang Uniformity (BU) [67]
5 Differential–Linear Cryptanalysis Linear Structures (LS) [65, 66]; Differential Linear Connectivity Table (DLCT) [69]; Differential Linear Uniformity (DLU) [70]
6 Algebraic Cryptanalysis Algebraic Degree (AD) [71]; Interpolation Polynomial (IP)
7 Side-Channel Analysis DPA-SNR [72]; Transparency Order (TO) [73]
Table 4: Theoretical Bounds on the Cryptographic Properties of S-boxes
S# Properties Bounds